OpenStudy (anonymous):

Graph the following function, state the domain, range, and asymptote. y=-2^x+3

OpenStudy (anonymous):

Lets see your function is a logarithmic function is it \[2^{x+3}\] or is it \[2^{x} +3\] I just want to be sure, there are some cases of confusion before when I did these types of questions =))).

OpenStudy (anonymous):

woops actually its a exponential function...lol

OpenStudy (anonymous):

well you forgot the negative and its the second one

OpenStudy (anonymous):

ok... so the domain is \[x \in \mathbb{N} \] Its an element of all real numbers and \[y \in \mathbb{N} / y \le 3\] I think ... heres what I think the graph looks like...

OpenStudy (anonymous):